OpenStudy (anonymous):

An eighth-grade student claims she can prove that subtraction of integers is commutative. She points out that if a and b are integers, then a – b = a + -b. Since addition is commutative, so is subtraction. What is your response?

jimthompson5910 (jim_thompson5910):

Subtraction is not commutative. We can see this with a counterexample. Let a = 2 and b = 3 So a - b = 2-3 = -1 but... b - a = 3-2 = 1 So because -1 does not equal 1, this tells us that a - b does not equal b - a So subtraction is not commutative.

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np